B is 16 years younger than Aand one half B’s age is equal to one third of A’s age ..how old are they ??

 

Grade:10

4 Answers

Ridhiksha hemanth mulya
8 years ago
????
V. Harshit -
8 years ago
A is 32+16=48 years old and B is 32 years old ,....
x/2=(x+16)/3
Abishek arun
8 years ago
B is 32 years old
A is 48 years old
A = B + 16
B/2 = A/3
3B = 2A
1.5B = 1A
1.5 B = 1 B + 16
0.5 B = 16
B = 32
V. Harshit -
8 years ago
if u think i am coorect please approve